OpenStudy (anonymous):

Express in lowest terms. 27a^3-1/9a^2-6a+1

OpenStudy (anonymous):

9a^2+3a+1/3a_1 B 1/2 c; 3a-1

OpenStudy (anonymous):

A;9a^2+3a+1/3a_1 B 1/2 c; 3a-1

OpenStudy (anonymous):

@sue101 help

OpenStudy (anonymous):

\[27a^3-1/9a^2-6a+1\]

OpenStudy (anonymous):

i need help @campbell_st

OpenStudy (campbell_st):

quick question is it \[\frac{27a^3 -1}{9a^2 - 6a + 1}\]

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

ok... that makes it a lot easier the numerator is the difference of 2 cubes a^3 - b^3 = (a -b)(a^2 +ab + b^2) a = 3a b = 1 the denominator is a perfect square can you rewrite both expressions in a factored form..?

OpenStudy (anonymous):

c; 3a-1

OpenStudy (campbell_st):

nope re-write it first, then cancel the common factors...

OpenStudy (anonymous):

B 1/2

OpenStudy (anonymous):

B ;1/2

OpenStudy (campbell_st):

ummm unsure... is the denominator double the numerator...?

OpenStudy (anonymous):

okey 9a^2+3a+1/3a_1

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... how did you get that..?

OpenStudy (anonymous):

he denominator is a perfect square can you rewrite both expressions in a factored form..?

OpenStudy (campbell_st):

ok...so what does the factored form look like..?

OpenStudy (anonymous):

i think 9a^2+3a+1/3a_1

OpenStudy (anonymous):

@Yahoo!

OpenStudy (campbell_st):

lol... no the factored form before anthing is cancelled. Thats the skill.... you need to learn...

OpenStudy (anonymous):

okey thanks